how many mg of a metal containing 45% nickel must be combined with 6 mg of pure nickel to form an alloy containing 78% nickel?
Answer by htmentor(1343)

Let x = the number of mg of 45% nickel alloy
After mixing with 6 mg of pure nickel, total amount of alloy is x + 6
Therefore the mixture contains this amount of nickel:
0.45x + 6 = 0.78(x + 6)
Solve for x:
0.45x + 6 = 0.78x + 4.68
x = 1.32/0.33 -> x = 4
Ans: 4 mg
Check:
(0.45(4) + 6)/(4 + 6) = 7.8/10 = 0.78